Dilly Farm Supply is located in a small town in the rural west. Data regarding the stores operations follow: 18 • Sales areDilly Farm Supply is located in a small town in the rural west. Data regarding the stores operations follow: 20 • Sales are
Answer #1

18) Net income = (542000*75%)-40000-50000 = 316500

Retained earnings at the end of December = 227425+316500 = 543925

20) December purchase = (330000*70%)+(230000*70%*90%)-(330000*70%*90%) = 168000

Account payable = 168000

So answer is a) $168000